I am currently 5 times as my son in 6 years i will be three times as old as he will be then. what are our ages now?
Answer by lwsshak3(11628)

I am currently 5 times as old as my son. In 6 years i will be three times as old as he will be then. what are our ages now?
..
let x=son's current age
then 5x= father's current age
6 years hence:
x+6 = son's age
5x+6=father's age
5x+6=3(x+6)
5x+6=3x+18
2x=12
x=6
ans:
son's current age=6
father's current age=30